Write a recursive formula for the sequence and find the next three terms in the sequence.
Sequence: 2, 11, 38, 119, . . .

Answer by stanbon(75887)

Write a recursive formula for the sequence and find the next three terms in the sequence.
Sequence:
a(1) = 2
a(2) = 2+3^2 = 11
a(3) = 11+3^3 = 38
a(4) = 38+3^4 = 119
-------
a(n) = a(n-1)+ 3^n
===========================
Cheers,
Stan H.
===========================